Status Report

NASA Space Station On-Orbit Status 2 Apr 2004

By SpaceRef Editor
April 2, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 2 Apr 2004

All ISS systems continue to function nominally, except those noted previously or below.

Update on Noise:  Early this morning (~2:30am EST) the crew reported another occurrence of the noise heard on 11/26/03 (then described as a “whumping sound, similar to when a thin sheet of metal is shaken”), or at last very similar to it, both in sound and location.  There was no apparent systems anomaly associated with it.  Investigation continues. [One of the current working theories focuses on a fan in the SM/crew quarters as source, with the noise transmitted through the ductwork.]

FE Alexander Kaleri set up and configured, with ground support, the hardware for the Russian geophysical GFI-1 “Relaksatsiya” (“relaxation”) payload and conducted a calibration using the Moon.  Subsequently, the equipment was torn down and re-stowed.  [After reconfiguring the payload Laptop 3 for the activity, Alex unstowed six payload accessories kits and installed the ultraviolet camera (UFK) with spectrometer unit at Service Module (SM) window #9.  From 3:51 to 4:02am EST, he conducted the Fialka (“violet”) experiment that uses the (known) spectra of the Moon and stars for payload calibration.  He also recorded voice comments on the observed image on the LCD (liquid crystal display) screen.  Relaksatsiya deals with the study of the chemoluminescent chemical reactions and atmospheric light phenomena that occur during high-velocity interaction between the exhaust products from space vehicles and the atmosphere at orbital altitude and during the entry of space vehicles into the Earth’s upper atmosphere.]

The FE performed another daily inspection session on the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.

Due to BCAT-3 (Binary Colloid Alloy Test 3) sample processes progressing much faster than anticipated, an additional photo session of specimen #1 with new camera settings was added to Mike Foale’s schedule today.  BCAT photos for the first four sessions have been successfully verified, and the ground team is “extremely excited” by the early (proprietary) results.

After the Lab’s EXPRESS Rack 4 (ER4) was power-cycled by POIC (Payload Operations & Integration Center, Huntsville), the Science Officer activated the SAMS-II (space acceleration measurement system 2) control panel to allow the ground to continue troubleshooting the payload. [The problem is believed to be between a controller in the ICU (interim control unit) and the hard drives.  SAMS-II measures accelerations caused by vehicle, crew and equipment disturbances.  These transient vibratory forces occur in the frequency range 0.01 to 300 Hertz.  The SAMS ICU consists of one of the PCS (portable computer system) ThinkPad laptops, operating across the Ethernet network known as the Medium Rate Link (MRL) in a two-way communication mode similar to that used to network computers together between offices.]

Alex Kaleri terminated the current measurement session with the Molniya-SM/LSO equipment, running on SM window #3 since 3/28, and deactivated its hardware, powering off the EGE-1 laptop. [Objective of the unattended Molniya-SM session, similar to the French LSO experiment, was to record storm phenomena and other related events in the Earth’s equatorial regions.  The experiment is controlled from the French EGE-1 laptop, loaded with orbital sighting predictions using up-to-date NORAD tracking TLE (two-line element) “slug” provided by NASA.  Objective of LSO is to study rare optical phenomena occurring in the upper layers of Earth’s atmosphere, so-called “sprites” (i.e., puzzling glow phenomena observed above thunderstorm clouds).  LSO was originally part of Claudie Haigneré’s French “Andromeda” payload package of taxi mission 3S, which could not be performed during Increment 4 as planned due to an ISS flight attitude conflict.]

The Elektron O2 generator continues to operate nominally, now in 24-amp mode.  To replenish its water supply for electrolysis, Sasha refilled the KOV thermal loops’ EDV container with purified water from the BKO multifiltration/purification column unit and the GZhS air/liquid separator unit while checking for any air bubbles (and, if visible, estimating their number in the EDV). [This recently introduced new procedure, intended to prevent gas bubbles from getting into the BZh fluid unit, now requires that the water, coming from a CWC (contingency water container) via the BP pump unit, first passes through the BKO and GZhS before its introduction into the BZh.  Formerly, when tapping a CWC supply, this was not the case.]

Mike completed the regular task of transferring data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm.  Later, he transferred the accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleted them on the HRM.

The crew had their daily workouts on TVIS, RED exerciser and VELO cycle with load trainer.

Sasha Kaleri held his regular weekly IMS (inventory management system) tagup with ground specialists, discussing open issues concerning identification of equipment and storage locations for updating the IMS database. [Today’s topics dealt with location updates for the laptops involved in the recent laptop “musical chairs” and some HDDs (hard disk drives).]

The crew continued pre-loading cargo preparations and equipment transfers to Progress 13P for preliminary loading activities, guided by reference to the IMS.

Alex attended to the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities) and prepared the daily IMS “delta” file for updating the database, while Mike conducted the regular routine status checkup of payloads currently running autonomously in the Lab.

Using the U.S. CDMK (carbon dioxide monitoring kit), Sasha took the weekly CO2 partial pressure reading of the SM cabin air, for calldown to the ground (along with the battery status) for trending analyses, then restowed the CDMK at the SM Central Post.

Mike Foale was commended on yesterday’s “excellent job” on the robotics operations with the Canadian SSRMS (space station remote manipulator system), which were completed with no issues.  Today, the CDR again disconnected the UOP/DCP (utility outlet panel/display & control panel) power bypass cable.

At about 1:55pm, the crew had their weekly teleconference with the ISS Flight Director at MCC-H.

As a routine part of the Russian health maintenance program, TsUP/Moscow uplinked recommendations to Kaleri, as for all expeditions approaching their end of mission, for taking preventative cardiotropic medication (i.e., drugs intended as countermeasure against the possibility of heart muscle shrinkage in zero-G).

Yesterday’s inspection of the ISS Leak Patch Kit revealed that both foam applicators are leaking.  They were immediately triple-bagged as per instructions and stowed for return to Houston. [The silicone foam is assigned Toxicity level 1, and the crew was properly protected for the inspection.  Although the foam applicators remain available for use in an emergency situation to repair leaks to vacuum, they are considered degraded, and their effectiveness cannot be guaranteed.  The primary method for leak repair thus becomes the Duxseal, a heated paste substance that was freshly flown up on Soyuz 7S and is also part of the Leak Patch Kit.]

U.S. checkout of the TRRJ (thermal radiator rotary joint) software was completed. [Some suspected flight software errors in the FDIR (failure detection, isolation & recovery) algorithms will be corrected with a patch for the software (Vers. R4) used in the EXT MDMs (external computers).]

Yesterday’s upload modification of the U.S. GNC MDMs (guidance, navigation & control computers) software with a time-tag patch for its RGA (rate gyro assembly) data-acceptance function was completed successfully, after attitude control handover to RS thrusters. [GNC MDM data are being dumped today for the ground to verify correct working of the patch.]

For the proper pointing of the GFI-1 Relaksatsiya spectrometer (see above), station attitude, still in LVLH (local vertical/local horizontal), was commanded to experiment attitude at 3:35am, after control handover to the Russian segment (RS) at 3:25am.  Later, at 4:05am, the station was maneuvered to XPOP (x-axis perpendicular to orbit plane) attitude, where it will remain until 4/21 (after Soyuz 8S docking).

Today’s CEO (Crew Earth Observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Urumqui, China (ISS had a near-nadir pass in fair weather over this remote city in western China, situated on the southwestern edge of the Dzungarian Basin), Rio de Janeiro, Brazil (the time of this pass should have been just prior to the daily formation of cumulus clouds.  Looking slightly right of track for good views of this famous Brazilian coastal city), Dakar, Senegal (mainland Africa’s westernmost city is situated on Cape Verde, just left of track), Rome, Italy (fair weather was expected this pass for good views of the Eternal City, just left of track), Lower Amazon River Basin (taking advantage of this break in the weather over this region to acquire near-nadir mapping strips of this rapidly changing landscape), Caracas, Venezuela (the Venezuelan capital is a challenging target to see, situated just inland.  Looking just left of track as ISS approached the coast), London, England, Great Britain (observers are not sure how much longer the unusual blocking pattern for weather over northern Europe will persist, but ISS appeared to have another chance for London, just left of track this pass), and Grand Banks/Flemish Cap (an unusual patch of very clear conditions was noted today SE of Newfoundland.  If this persisted, ISS had good sun glint views for internal waves here to the right rear of its ground track).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of today, 1:35pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair not completed; to be tested ASAP — see above note).  SFOG slot#2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.9; ppO2 (mmHg) — 159.5; ppCO2 (mmHg) — 3.7;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.0.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 754.07; temperature (deg C) — 23.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 755.83; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 756.03; temperature (deg C) — 23.0; shell heater temp (deg C) — 23.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — 24.3
  • PMA-2:  Shell heater temp (deg C) — 14.4

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind mode, non solar-tracking, drag reduction-biased with 47 deg).
  • SM batteries:  Battery #8 is off-line; battery #2 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries:  Battery #6 is off-line (capacity restoration mode, ROM); battery #1 is in “Cycle” mode (PTAB suspect); all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Discharge mode; PCU-2 is in Discharge mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3926 kg (8655 lb) as of 3/26/04;  [SM(755) + FGB(2512) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Current SSC (Station Support Computer) Laptop Status:

  • 7 SSC ThinkPad A31p laptops in use (1 File server, 5 OOC, 1 crew personal)
  • 1 ThinkPad A31p in use (as CPSD/crew personal support drive, not on network)
  • 1 ThinkPad A31p at SM Central Post (on loan to RS & converted to LT1).
  • 6 SSC ThinkPad 760XD laptops in use (4 SSC Clients, 1 SSC Router, 1 OCA Router)
  • 1 ThinkPad 760XD stowed (former SSC2, reboot error)
  • 1 ThinkPad 760XD failed (former OCA Router)

  • Robotics:
  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:29am EST [= epoch]):

  • Mean altitude — 365.4 km
  • Apogee — 372.7km
  • Perigee — 358.2 km
  • Period — 91.85 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.001073
  • Solar Beta angle — 27.8 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.68
  • Mean altitude loss last 24 hours — 130 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30644

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.